Please use a calculator for the following problems and please show work

1) Brandie wants to deposit $10,000 a year for 30 years. She then wants to retire for 30 years using all the monies available equally. How much can she withdraw for the 30 years assuming an interest rate at 4%

2) Julie currently has on hand $30,000 that will be deposited in a savings account until it accumulates to $50,000. How long will it take Julie to accumulate the $50,000, assuming 4% compounding on an annual basis

3) What would you pay for a $200,000 bond that matures in 10 years and pays 8% interest (stated or coupon rate0 a year if you wanted to earn a yield of 6% (market rate)
Answer #1

ANSWER

1)

Present value of savings = 10000*(1-1/1.04^30)/4%=172,920.33

Future value of savings at the end of 30 years = 172,920.33*1.04^30 =560,849.37

Let withdraws be x

560,849.37= x*(1-1/1.04^30)/4%=

X= 32,434.04

she withdraw for the 30 years = $32,434.04

2)

50,000 =30,000 *1.04^n

N = 13.02 years

Time = 13 years approx

3)

Coupon = 8%*200000 = 16000

Price = 16000*(1-1/1.06^10)/6% + 200000/1.06^10=229,440.35

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************